2 8 25 25 The above table has the demand and supply schedules for money. Real GDP increase s and, as a result, the demand for money t increases by $0.2 trlion at each level of the nominal A. 10 percent B. 7 percent. OC. 2 percent. D. 5 percent. C. E, 3 percent. 6. The GDP price index in the United States in 2002 was about 85, and real GDP in 2002 was 31 2.9 trillion (2009 dollars). x in the United States in 2010 was about 99, and real GDP in 2010 was 514.8 trillion (2009 dollars). Calculate nominal GDP in 2002 and in 2010 an d the percentage increase in nominal GDP between 2002 and 2010 >>>Answer below to 1 decimal place. Nominal GDP in 2002 is $ Nominal GDP in 2010 is $ The percentage increase in the nominal GDP between 2002 and 2010 is the nominal interest rate is 10 percent, the inflation rate is 6 percent, and the tax rate on interest income is 25 percent, what is the after - lax real interest rate? A. 3.5 percent trillion. 17. If B. 4.0 percent O c. 1.5 percent D. 3.0 percent E. 6.0 percent
Answer #1

Ans . 7% - The eql interest rate becomes 7%

16) Nominal GDP (2002) = 85 × 12.9 ÷ 100 = 10.9

Nominal GDP ( 2010) = 99 × 14.8 ÷ 100 = 14.6

% Increase in Nominal GDP = 33.9%

17) 3% - The after tax Real interest rate would be 3%

Best of Luck !! Thank You !!

A Thumsup would be Highly Appreciated !!

!!